### v1

# Economic thermodynamics and inflation

Abstract
This study presents a computational and theoretical framework inspired by thermodynamic principles to analyze the dynamics of economic inflation within adiabatic and non-adiabatic systems. In a framework referred to as developmental symmetry, inflation is formulated as a scalar field evolving through continuity equations, drawing an analogy with the Raychaudhuri equation in gravitational dynamics. The results show that adiabatic systems fail to reach equilibrium, while non-adiabatic systems can evolve toward stable states over time. The model successfully reproduces observed inflationary regimes-from hyperinflation to stable low-inflation phases-with characteristic transition periods of about a decade. These results indicate that production continuity and controlled monetary flow are crucial for achieving stability in complex economic systems, linking thermodynamic balance to macroeconomic equilibrium.
Inflation, Raychaudhuri equation, Economics Equilibrium, Thermodynamics
I Introduction
Inflation is characterized by a rapid decline in the purchasing power of money and a seeming monotonous upward movement of prices. Inflation has occurred frequently throughout historyR1 , and some nations are still dealing with its threat. So-called hyper-inflation, in particular, is a type of monetary system failure that is detrimental to society as a whole. Hyperinflation models are particularly useful for highlighting how inflation indicates poor ”states of natural system” in the economy. Cases of such a system are characterized by wars, social regime shifts, state bankruptcies, etc. As can be seen in the analyzed situationsR2 ; R3 ; R4 ; R5 ; R6 ; R7 ; R8 ; R9 ; R10 , these elements, along with the influence of pure economic variables like expectations, money demand, velocity of circulation, and quantity of money, result in a rise in the consumer price index that is greater than exponential. This behavior therefore has a detrimental impact on the social network, creating awkward circumstances. Since an exact scientific definition of inflation has not yet been established, there is currently no method available to detect hyperinflation in its early stages, which could help prevent the tragic event. The literature-available model for hyperinflation is based on a nonlinear feedback indicated by a power-law exponent with a value greater than 0. In such a method, it can be mention that the consumer price index displays a finite temporal singularity of the type, $\frac{1}{(t_{c}-t)^{\frac{1-\alpha}{\alpha}}}$, enabling the identification of a critical moment $t_{c}$ at which the economy would collapse. Numerous situations where this model has been used successfully can be observed in studies R4 ; R5 ; R6 ; R7 ; R8 ; R9 ; R10 . Real economic activities are harmed when inflation rises over modest levels. For example, because tax revenues are effectively received over a longer period after the announcement that fixes them, they have an impact on government revenueR11 ; R12 . Since it is difficult to tell whether a price increase is due to a relative price shift or is a result of general inflation, perceiving relative price changes becomes more challenging, which is called the Lucas problemR13 . If the adjustment process varies for different types of goods, it results in an inaccurate allocation of resourcesR14 and leads to wasteful changes in relative pricesR15 . A currency’s qualities such as a store of value, a medium of exchange, and a unit of account are all impacted by inflation. As a result, the level of disturbance increases as inflation increases.
The thermodynamic analysis of an economic system (the first and second law, thermodynamic equilibrium and processes) and analysis with methods in statistical physics can be seen in the literatureR16 ; R17 ; R18 ; R19 ; R20 ; R21 ; R22 ; R23 ; R24 ; R25 ; R26 ; R27 ; R28 ; R29 ; R30 ; R31 . In general, it can be said that the analysis of both economic and social systems, with their appropriate definitions in physical sciences, is a form of ”developmental symmetry”. In particular, by drawing an analogy with its application to physics, it has been demonstrated in work R30 ; R31 that the thermodynamics of markets can be constructed as a phenomenological theory.
In this study, it is shown that Raychaudhuri equation R32 ; R33 ; R34 ; R35 can describe an economic system without going into deep mathematical proofs. The solutions of the equation for adiabatic and non-adiabatic processes that occur in the energy conservation equation, which are considered in the thermodynamic framework, are investigated. At this point, it has been observed that the adiabatic process in which hyperinflation emerged can be eliminated by certain economic processes and monetary policies directing the system. In this direction, we consider the economic system as a field that can be defined as a scalar field $\phi$. The existence of a minimum inflation in this field is a necessity even at the beginning. The Raychaudhuri equation appears as an equation describing such an economic system that can bend and expand continuously. When we say bending or curvature of the field, we mean whether the unit (point A in Figure 1) with kinetic energy (monetary purchasing power) can reach a point with potential energy (movable-real estate, valuables) within the field.
II The Raychaudhuri equation and inflation
In general, inflation is a concept used to explain the increase in average prices in the economy. On the other hand, this also means that money loses its value (Reserve Bank of New Zealand R36 ). In the mathematical constructs of economics, rather than theoretical proof, economic conceptual phenomena are structured within the framework of theory. With these intuitive and reasonable assumptions, economic systems are followed and analyzed through the lens of theory. We will mathematically restructure inflation dynamics, which expresses the constant increase in the price of movable and immovable goodsR31 , not in a part of the economic system but in its general structure, with a scalar field. Similar to cosmological inflation, we assume that scalar fields have a role in causing inflation in economy. In this context, we refer to ”developmental symmetry” as the property of a phenomenon in the observable universe having a structure that is similar to or symmetrical with a phenomenon in the social structure. Considering the fact that inflation takes place in the entire region of a real economics system, within the framework of thermodynamic boundary we assume that the inflation occurs and in this way is distributed throughout all point of the homogeneous structure (economics system).
Let the purchasing power be represented by point $A$ in the field and the price of the good by point $B$. The distance between both points is scaled by $x$. However, we consider the condition of availability of $B$ in an economic system (no deflation). It is worth emphasizing that when the acceleration term is $\ddot{x}>0$ it states the field is in the inflation phase and at $\ddot{x}<0$ it is in the deflation phase. Here, it should be apprehensible that an increase in the x scale under the $\ddot{x}>0$ condition indicates a decrease in purchasing power or shopping, thus indicating inflation. In the literature, the relative rate of change in the price of the good is given by the equality $C=\frac{\dot{p}}{p}$R37 ; R38 ; R39 ; R40 . According to the method we have discussed in this study, we give another definition of the inflation rate of change. For this purpose, one can define inflation rate $C=\frac{\dot{x}}{x}$ that shows how an economic system or inflation field has evolved. In such a system, the emergence of a developed economic equilibrium depends on the system’s conversion of $B$, which is formed by production, as added value to the purchasing power at point $A$. In an inflationary system, the most important way of converting movable-immovable goods (including technological production) into monetary resources is exportation. Another known tool is the resources transferred to the system (country). Such systems describe non-adiabatic systems that are a real phenomenon of the global system. On the other hand, it can be easily seen that the availability of purchasing power of $A$ to $B$ decreases continuously since the rate of $C$ tends to decrease over time under the inflationary condition. This case, for example, can be observed on the $x\simeq t^{h}$, where a decrease in $C$ is observed at the rate $\frac{h}{t}$ over time when $h>1$. Therefore, it is seen that the economic system has a dynamic structure with $x$ scaling.
Figure 1: Representation of any subsystem chosen in the $E$ economic system. Points AA and $B$ in the subsystem are scaled by $x$. The AA unit of density (a buyer) defined in $A$ is represented by kinetic energy (purchasing power). Point $B$ is the potential energy density representing movable-immovable goods or various assets. During inflation, there exists pressure caused by the scalar field towards the outside of the system $E$, and the system always has a dynamic and expanding feature. While it shows continuous growth on the $x$ scale in the expanding field, it should be controlled with fiscal policies suitable for the system for purchasing power and goods price balance.
The energy density and pressure of this homogeny and dynamic system are given as followsR41 ; R42 ,
$\rho=\frac{\dot{\varphi}^{2}}{2}+V(\varphi),\,\,\,\,\,\,\,\,\,\,p=\frac{\dot{\varphi}^{2}}{2}-V(\varphi)$ (1)
Here, the kinetic energy is the purchasing power at the $A$ point, and the $V(\varphi)$ potential energy represents all the movable and immovable goods at the $B$ point. It is clear that the decline in $C$ reduces the accessibility and so defines the relationship between purchasing power and goods (price). In this regard, the quantities $A$ and $B$ are important objects of the system for our discussion. It is useful to note that inflation occurs when $\frac{\dot{\varphi}^{2}}{2}<V(\varphi)$. The economic case is at a desired level (economics equilibrium) when $\frac{\dot{\varphi}^{2}}{2}\simeq V(\varphi)$. The first ($\frac{\dot{\varphi}^{2}}{2}<V(\varphi)$) is when the price of goods is high but purchasing power is low. This represents a phase in which the $x$ scale continuously increases very rapidly. The second one indicates the existence of an economic equilibrium. In the second case, there is a slow increase in the $x$ scale, which is a necessary condition for a developed country’s economy. Therefore, the first case produces negative pressure corresponding to inflation (expanding of the field), and the second one gets positive pressure corresponding to purchasing power (deceleration of the field). On the other hand, we proceed by defining a modulating parameter, $\gamma=\frac{p_{\varphi}}{\rho_{\varphi}}$, which also denotes a state of the equation. While it is an inflation case with $\gamma<0$, the positive definition of $\gamma$ shows a non-inflation case (decelerating case in the between points).
We start with the Raychaudhuri equationR32 ; R33 ; R34 ; R35 , which describes the motion of localized units (densities) at points AA and $B$ defined in the field, and describes the motion between two points. For a homogeneous distributed dynamic system, this equation can be written as follows:
$\frac{d\theta}{d\tau}=-\frac{1}{3}\theta^{2}-4\pi(\rho_{\varphi}+3p_{\varphi}).$ (2)
where $\theta=\frac{\dot{x}}{x}$. Therefore, it can be written the equation,
$\frac{-3\ddot{x}}{x}=4\pi(\rho_{\varphi}+3p_{\varphi}).$ (3)
However, to achieve the continuity equation of the perfect fluid we make use of the first law of thermodynamics,
$\delta Q=dE+pdV=TdS,$ (4)
where $E$ is internal energy of subsystem (amount of money in the system), $V$ is the volume of system (the amount of goods in system), $p$ is the pressure of system (the indicator of purchasing power in the system) and $\delta Q$ is heat follow (money transfer that occurred between the subsystems or systems without buying and selling goods). Hence, since a macro-level observer will observe the energy from these quantities as a density dispersed in the volume in the subsystem, under the approximation $dE\simeq dm\simeq Vd\rho+\rho dV$, the first law is read as follows,
$\dot{\rho}+3C(\rho_{\varphi}+p_{\varphi})=\frac{\delta Q}{Vdt},$ (5)
This is the conservation equation of the economic system, which states that the quantities of the system are conserved during the economic cycle with time. Two solutions of equation (5) can be investigated for adiabatic ($\delta Q\simeq 0$) and non-adiabatic ($\delta Q\neq 0$) processes. The adiabatic case where there is no money transfer at the borders of the system means that there is no money transfer economically to a country without buying or selling goods. The other non-adiabatic process means there is a flow of money to or from country borders. In the next two sections, we will discuss the solutions of these two processes from an economic point of view.
II.1 Adiabatic process
The continuity equation for the adiabatic process takes the following form,
$\dot{\rho}+3C(\rho_{\varphi}+p_{\varphi})=0,$ (6)
and whose solution is as follows
$\rho_{\varphi}=\rho_{0}a^{-3(1+\gamma)}.$ (7)
Here, $\rho_{0}$ is an integral constant. Then, the eq. (5) takes the form,
$3\ddot{x}=4\pi\rho_{0}x^{-3(1+\gamma)+1}(1+3\gamma),$ (8)
So, we proceed by finding some solutions for the eq. (9) to investigate inflation. Using the modulating parameter,$w$, we can write $\dot{\varphi}^{2}(1-\gamma)=2(1+\gamma)V(\varphi)$. This equation tells us that there exists a boundary in an economic system at the point $\gamma\neq 1$ because the case $V(\varphi)=0$ representing movable or immovable property is not considered in an economic system. But, it is possible in the case $\gamma\sim-1$, which corresponds to a hyperinflation. This is a case where the monetary purchasing power represented by kinetic energy, $\dot{\varphi}^{2}$, is almost nonexistent or capital becomes almost worthless. Therefore, the solutions close to $\gamma\sim-1$ produce exponential hyperinflation solutions,
$x\simeq\exp(\sqrt{\frac{8\pi\rho_{0}}{3}}t)C_{1}+\exp(-\sqrt{\frac{8\pi\rho_{0}}{3}}t)C_{2},$ (9)
where $C_{1}$ and $C_{2}$ are integral constants. The first solution, which shows the increase in the distance scale between the purchasing power and the price of the good, is considered. In other words, an increase in scale $x$ means that there is a decreasing effect on purchasing power due to $\gamma\sim-1$ ($\frac{\dot{\varphi}^{2}}{2}\ll V(\varphi)$). Therefore, we can set $C_{2}$ to zero, where the solution $x\simeq\exp(\sqrt{\frac{8\pi\rho_{0}}{3}}t)C_{1}$ survives. This appears in Figure 1 with $C_{1}=1$.
Figure 2: The relationship between the distance scale and time is given. An exponential increase in the $x$ scale corresponds to hyperinflation. In the $\frac{\dot{\varphi}^{2}}{2}\ll V(\varphi)$ expression, purchasing power (kinetic energy) becomes almost negligible, and the price increase is caused by high potential (movable and immovable property, various assets).
The high potential indicates the high price of material resources that have not yet been converted into monetary (kinetic energy) resources. An exponential increase in the $x$ distance scale indicates a decrease in purchasing power. This unstable situation can be eliminated to some extent by reducing the demand for the purchase of goods, which leads to a decay of the potential, that is, to a decrease in price. However, in such an adiabatic system (that is, there is no monetary input and output to the system), the potential energy (price) can be reduced by increasing production, or, as we emphasized, the price of resources can be reduced by decreasing demand. An important point to be emphasized here is that forcing the system into an imaginary non-adiabatic state is not a solution. Because, even if a heat transfer (monetary resources that can come from outside the country, for example, transferring money to the system with interest) increases the kinetic energy of the system in the short run, it will lead to heavier inflation in the long run. This can be seen in Figure 2 from the peaking of the graph with increasing time. In addition, non-real temporary monetary resource inflow for the system forces the inflation field to a heterogeneous structure in real terms. In the long run, increasing the virtual kinetic energy in the system will lead to the draining of the resources of the system with interest and lead to more imbalance in the densities within the area. Because monetary resources will lead to the gain of one side and more exploitation of the other side, which will result in the emergence of an inhomogeneous area. In real terms, the purchasing power of density units (persons) whose kinetic energy decreases more will weaken and high inflation will continue. These are some conditions that apply to the adiabatic system. Unit densities in adiabatic systems appear to be a more realistic way to reduce demand (making economic savings) for goods, reduce inflation, and bring the economy back to reasonable equilibrium. In this case, the seller will have to reduce the price to convert his movable-immovable or various assets into money, which is one of the most important factors in reducing inflation. On the other hand, by revealing the natural resources in the system, by increasing the resources, the price can be lowered and economic stability can be achieved. Natural resources will lead to the degradation of potential and an increase in monetary (kinetic energy) valuation. Indecision will continue, as wasting resources is contrary to the approach and policy of buying fewer goods. It should be noted here that being satisfied with the goods in hand will either be implemented by the units themselves or economic policies will force the system to do so. The change in the rate of inflation in this phase is equal to $C=\rho_{0}$. If we fix the value of $\rho_{0}$ at a very large value, inflation becomes continuous with an approximately unchanged value of $C$. Therefore, within the framework of reasonable policies, this rate of change parameter should take a time-dependentR42 form that will undergo time evolution. In this direction, in developed economic systems, the system should reach the equilibrium point after inflation. This equilibrium point can be thought to be possible if the purchasing power is equal to the potential, $\frac{\dot{\varphi}^{2}}{2}\simeq V(\varphi)$, corresponding to the point $\gamma\simeq 0$. But, the solution of the eq. (10) around at this point, $\gamma\simeq 0$, is an empty set. Therefore, although $K=V$ for adiabatic systems, as a result of reasonable policies (example of reducing the demand for goods by saving), we obtain a linear solution as follows at the point $\gamma\simeq-\frac{1}{3}$ with the falling pressure of the system.
$x\sim C_{1}+tC_{2},$ (10)
with real part ($C_{1}=0$), we have $x\sim tC_{2}$. The ratio parameter is in the form $C=\frac{1}{t}$, which exhibits a decreasing behavior with time. As a result, fiscal policies that would require a long time under the constraint of the adiabatic system (economy or fiscal restraint policies) may result in lowering inflation. If $t$ is taken as years, it may take more than 10 years for inflation to be expressed in single digits. As we will see in the next chapter, a reasonable economic equilibrium point may emerge at $Kinetic(K)=Potential(V)$, and a decrease in inflation can be observed in less than 10 years with inter-system trade, except for fiscal constraints or the position of unit densities to settle for less goods.
II.2 Non-adiabatic process
We shall use the entropic force definition to reveal how the term on the right-hand side of eq. (7). The entropic force is defined as follows,
$F_{ent.}=-\frac{TdS}{dx},$ (11)
where for a non-adiabatic process we have $TdS=\delta Q\neq 0$, so the eq. (7) can be written as follows,
$\frac{d\rho}{\rho}+3(1+\gamma)\frac{dx}{x}=-\frac{F_{ent.}dx}{E}.$ (12)
Herein, due to $E\simeq\rho V\simeq F_{ent.}x$, where recall that $\rho=kinetic+potential$, we can write
$\frac{d\rho}{\rho}+3(1+\gamma)\frac{dx}{x}=-\frac{dx}{x}.$ (13)
The solution of this equation is as follows,
$\rho=\rho_{0}x^{-3(1+\gamma)-1},$ (14)
with an integral constant $\rho_{0}$. Thus, the eq. (5) takes the form,
$3\ddot{x}=4\pi\rho_{0}(1+3\gamma)x^{-3(1+\gamma)}.$ (15)
Hyperinflation solution $\gamma\simeq-1$ is
$x=\frac{4}{3}\pi\rho_{0}t^{2}+C_{1}+C_{2}t,$ (16)
where we set $C_{1}=C_{2}=0$, so, an increase in $x$ is parabolic, indicating very low economic purchasing power.
Figure 3: Figure showing the increase of the x scale in the (0-8) time interval.
The proportional parameter of this type of inflationary field behaves as $C=\frac{2}{t}$. Even if the inflation parameter $C$, which starts in a zero time instance, returns to its natural appearance after 20 years, it reveals a deep deformation situation that ruins the field due to the size of the $x$ scale. Therefore, in such an inflationary field, it is necessary to intervene in the system with reasonable fiscal policies and to pass the inflationary phase to a new phase. In this way, a stable course can be followed by keeping the economy under control without any deformation. Therefore, when we apply the $K=V$ or $w\simeq 0$ approximation that we discussed for economic equilibrium to the equation, we get the following equation.
$x=\pm(\frac{-4\pi\rho_{0}+3t^{2}C_{1}^{2}C_{2}+6tC_{1}^{2}C_{2}+3C_{1}^{2}C_{2}^{2}}{3C_{1}})^{\frac{1}{2}}.$ (17)
Note that the solutions (17) and (18) converge at about $t=4$. We can manipulate the value of $C_{1}$ based on any value of $C_{2}$. For example, the value graph at $C_{1}=0.4$ for $C_{2}=1$ in the wide selected range of x=(415)x=(4-15) is given in Figure 1. If $C_{2}=2$ is chosen, $C_{1}=0.34$ produces the same graph. This means that inflation entered a declining phase (from $t\sim 4.1$ to about 11) and progressed in a developed economy phase after $t\sim 11$.
Figure 4: The figure shows the behavior of the x scale in the time interval (4-15). Here, $C_{1}=0.4$ is chosen.
We can again read the change in the inflation rate from the parameter $C=\frac{\dot{x}}{x}$. If we consider $t$ in terms of years, the economy with the traces of hyperinflation between $t=4.1$ and $t=5$, and a rate that decreases to 10 percent in approximately 7 years between $t=4.1$ and $t\simeq 11$ is obtained. After the approximate value of $t\simeq 11$, a single-digit inflation value may appear. This indicates the economic equilibrium represented by $K=V$ for the economic system. Let’s note that it takes about 10 years for it to reach equilibrium from a state of almost no purchasing power during the hyperinflation period and that it takes place with the external heat transfer to the system (occurring money transfer without buying- selling cases). Let’s remember that the adiabatic process did not produce the $K=V$ solution. In such processes, we emphasized that inflation can be avoided by reducing the demand for purchasing goods.
Country Inflation period
Bolivia 1970-1985
Peru 1970-1990
Israel 1970-1985
Brazil 1970-1994
Nicaragua 1985-1991
Hungary 1946 Jan-1946 Jul
Germany 1922 Nov-1923 Oct
Table 1: Inflation periods in certain year intervals of some countriesR3 .
With the decrease in demand for the density (energy, that is, money) units that describe the system, the kinetic energy, that is, the purchasing power will increase over time. Another possible situation is the realization of the political intervention in the system to reduce the demand for the purchase of goods. Since this policy will reduce purchasing demand over time, the consumer’s decreasing list of needs will increase kinetic energy over time, thus increasing purchasing power. However, in the non-adiabatic process, external monetary investments in the system or the existence of rich resources (potential assets that can be converted into money, that is, kinetic energy) in the field of the system will push the process to an economic balance over time, which our solutions in this topic have shown. Table 1 shows some countries experiencing hyperinflationary periods. When considered yearly, it is observed that there exists an inflation process of about 10 years. This periodic process is in harmony with our calculations. On the other hand, under the condition of applying reasonable financial policies, the economic system in Turkey for example, which started in 2017 and evolved into a hyperinflationary process, may be expected to complete its natural course after 7 years and decrease to a single-digit after 11 years (around 2028 date).
III Closing remarks
There exist many symmetries in the universe (e.g. spatial and temporal translation symmetries, parity, and charge symmetries, etc.). Especially in physical science, symetry describes that when any observable quantity is considered, some properties of the system do not change under certain transformations (for example, Noether symmetries). In other words, symmetry tells us that the internal property of a physical system related to physics or mathematics does not change with the change of some factors. In social science, which is also a part of the universe, evaluating the welfare of society, the interaction of the units that make up the social system (including political interactions) and its development in terms of its compatibility with the laws of theoretical physics, can offer a realistic perspective. On the other hand, the physical development of a plant or human, the growth of the circle of ideas, or the expansion of the intellectual sphere of influence (just as the expansion of the universe) refer to observable symmetries. In this study, which we handled mathematically according to this point of view, which we call ”developmental symmetry” in the universe, we discuss inflation as an ever-expanding system (similar to the dynamic structure in developed economies) with its scalar field representation. In this direction, we investigate the economics of a system from an inflationary perspective within the framework of the Raychaudhuri equation. We have obtained a continuity equation based on the first law of thermodynamics. For the adiabatic and non-adiabatic processes of this equation, we obtain some solutions corresponding to the inflationary case for the system. As can be seen in Figure 1, the main point we consider in the study is the purchasing power, which is described by point $A$, and the purchased goods, which is identified by point $B$. It should be noted that in a developed economic system under consideration, the $x$-scale can never reach a value of zero. Finding an $x$ value in developed economies is important for the continuity and development of the economic system. Therefore, it is inevitable for a dynamic system (developed economic system) to have a constant value ($\frac{\dot{\phi}^{2}}{2}\simeq V(\phi)$) of purchasing power and price of goods for an exchange of approximately $x$ value to take place. An increase in the $x$ scale means that the availability of the good decreases, that is, the purchasing power ($\frac{\dot{\phi}^{2}}{2}$) decreases or the price of the good $V(\phi)$ is high. We represented this situation with $\frac{\dot{\phi}^{2}}{2}\ll V(\phi)$, therefore, solutions in which hyperinflation occurs in the case of $\frac{\dot{\phi}^{2}}{2}\ll V(\phi)$. As emphasized, the equilibrium point is the case of $\frac{\dot{\phi}^{2}}{2}\simeq V(\phi)$, which corresponds to a deflationary area that reduces the rate of hyperinflation. Since there is no energy (monetary) flow with external systems for an adiabatic economic system, it has been revealed that it can get rid of inflation in more than 10 years after it evolves into an inflation area of $\gamma=-\frac{1}{3}$ with reasonable financial policies within the system. Also considering the duration of reasonable policies, it will probably take more than 20 years to get out of the inflationary field. For the more realistic non-adiabatic process (the system with energy-money flow), we obtained mathematical results that are quite compatible with the inflation processes seen in Table 1.
On the other hand, the main reason why some countries in Table 1 left the hyperinflation area in less than 11 years can be shown as the inflow of money into the country. For example, the cash flow from the international monetary fund to the country arising from production, natural resources, technological infrastructure and the sound economic policies created as a result can be shown. The reasons for the inflation that continues for more than 11 years on average may be wrong fiscal policies and management of the system (country) as an adiabatic system. In our current study, the 11-year average period of the inflation process is obtained within the framework of ”developmental symmetry” for non-adiabatic systems. After an average of 11 years of hyperinflation, the economic system is evolving into a stable state (single-digit percentages). As we have emphasized, one of the most effective ways to stop hyperinflation in such an economic system created identically (symmetrically) to mathematical tools is to ensure production in the system. Whether such a system is adiabatic or not (whether there is energy-money flow or not) may no longer be important. The use of natural resources with the right policies or the production of technological parameters that may meet the needs can lead to a reasonable inflation rate without the flow of money with interest to the inflation area where purchasing power drops to almost zero. In another separate study, it can be discussed whether inflation can be cut or not by finding the creation of the scalar field density in the field. Because density is a quantity that includes kinetics and potential due to $E\sim\rho$. Units of density created in the field by the political ground controlling the field represents the unity of purchasing power (kinetic energy, point $A$ in Figure 1) and movable immovable property (potential energy, point $B$ in Figure 1). Therefore, it is clear that this unity emerging everywhere in the field will have a decreasing effect on inflation.
IV Data Availability Statement
No Data associated in the manuscript