### v2

# Economic thermodynamics and inflation

Abstract
This study presents a computational and theoretical framework inspired by thermodynamic principles to analyze the dynamics of economic inflation within adiabatic and non-adiabatic systems. In a framework referred to as developmental symmetry, inflation is formulated as a scalar field evolving through continuity equations, drawing an analogy with the Raychaudhuri equation in gravitational dynamics. The results show that adiabatic systems fail to reach equilibrium, while non-adiabatic systems can evolve toward stable states over time. The model successfully reproduces observed inflationary regimes-from hyperinflation to stable low-inflation phases-with characteristic transition periods of about a decade. These results indicate that production continuity and controlled monetary flow are crucial for achieving stability in complex economic systems, linking thermodynamic balance to macroeconomic equilibrium.
Inflation, Raychaudhuri equation, Economics Equilibrium, Thermodynamics
I Introduction
Inflation is characterized by a rapid decline in the purchasing power of money and a seeming monotonous upward movement of prices. Inflation has occurred frequently throughout historyR1 , and some nations are still dealing with its threat. So-called hyper-inflation, in particular, is a type of monetary system failure that is detrimental to society as a whole. Hyperinflation models are particularly useful for highlighting how inflation indicates poor ”states of natural system” in the economy. Cases of such a system are characterized by wars, social regime shifts, state bankruptcies, etc. As can be seen in the analyzed situationsR2 ; R3 ; R4 ; R5 ; R6 ; R7 ; R8 ; R9 ; R10 , these elements, along with the influence of pure economic variables like expectations, money demand, velocity of circulation, and quantity of money, result in a rise in the consumer price index that is greater than exponential. This behavior therefore has a detrimental impact on the social network, creating awkward circumstances. Since an exact scientific definition of inflation has not yet been established, there is currently no method available to detect hyperinflation in its early stages, which could help prevent the tragic event. The literature-available model for hyperinflation is based on a nonlinear feedback indicated by a power-law exponent with a value greater than 0. In such a method, it can be mention that the consumer price index displays a finite temporal singularity of the type, $\frac{1}{(t_{c}-t)^{\frac{1-\alpha}{\alpha}}}$, enabling the identification of a critical moment $t_{c}$ at which the economy would collapse. Numerous situations where this model has been used successfully can be observed in studies R4 ; R5 ; R6 ; R7 ; R8 ; R9 ; R10 . Real economic activities are harmed when inflation rises over modest levels. For example, because tax revenues are effectively received over a longer period after the announcement that fixes them, they have an impact on government revenueR11 ; R12 . Since it is difficult to tell whether a price increase is due to a relative price shift or is a result of general inflation, perceiving relative price changes becomes more challenging, which is called the Lucas problemR13 . If the adjustment process varies for different types of goods, it results in an inaccurate allocation of resourcesR14 and leads to wasteful changes in relative pricesR15 . A currency’s qualities such as a store of value, a medium of exchange, and a unit of account are all impacted by inflation. As a result, the level of disturbance increases as inflation increases.
The thermodynamic analysis of an economic system (the first and second law, thermodynamic equilibrium and processes) and analysis with methods in statistical physics can be seen in the literatureR16 ; R17 ; R18 ; R19 ; R20 ; R21 ; R22 ; R23 ; R24 ; R25 ; R26 ; R27 ; R28 ; R29 ; R30 ; R31 . In general, it can be said that the analysis of both economic and social systems, with their appropriate definitions in physical sciences, is a form of ”developmental symmetry”. In particular, by drawing an analogy with its application to physics, it has been demonstrated in work R30 ; R31 that the thermodynamics of markets can be constructed as a phenomenological theory.
In this study, it is shown that Raychaudhuri equation R32 ; R33 ; R34 ; R35 can describe an economic system without going into deep mathematical proofs. The solutions of the equation for adiabatic and non-adiabatic processes that occur in the energy conservation equation, which are considered in the thermodynamic framework, are investigated. At this point, it has been observed that the adiabatic process in which hyperinflation emerged can be eliminated by certain economic processes and monetary policies directing the system. In this direction, we consider the economic system as a field that can be defined as a scalar field $\phi$. The existence of a minimum inflation in this field is a necessity even at the beginning. The Raychaudhuri equation appears as an equation describing such an economic system that can bend and expand continuously. When we say bending or curvature of the field, we mean whether the unit (point A in Figure 1) with kinetic energy (monetary purchasing power) can reach a point with potential energy (movable-real estate, valuables) within the field. In this context, the present study aims to formulate an analytical framework in which macroeconomic behavior can be interpreted through the geometrical properties of the economic field. This approach provides a novel perspective for understanding inflationary dynamics and market evolution in terms of curvature and expansion processes. Thus, the proposed analogy offers a new interpretive framework connecting economic stability and inflationary behavior with field dynamics, suggesting that economic equilibria may be viewed as geometrical configurations within this scalar field. Therefore, this analogy provides a foundation for modeling economic evolution as a geometric process, allowing inflationary behavior and market expansion to be understood within a unified thermodynamic and relativistic framework.
II The Raychaudhuri Equation and Inflation
In general, inflation is a concept used to explain the continuous increase in the average price level within an economy. At the same time, this increase implies a decline in the value of money (Reserve Bank of New Zealand R36 ). In the mathematical constructs of economics, rather than relying solely on theoretical proofs, conceptual economic phenomena are structured within the framework of general system theory.
It should be noted that the present work does not aim to establish a strict physical equivalence between cosmological and economic variables. Instead, the analogy developed here is used phenomenologically to describe the qualitative behavior of economic expansion and contraction processes through the lens of dynamic systems.
With these intuitive and reasonable assumptions, economic systems are examined as evolving entities that can be analyzed using methods inspired by physical sciences. We will mathematically restructure the dynamics of inflation, defined as the persistent rise in the prices of movable and immovable goods R31 , not as a process limited to a subsystem, but as an intrinsic property of the general structure of the economy, represented within a continuous field-like framework.
Here, the economic variable $x(t)$ represents the relative distance between purchasing power and the price level, analogous to a system-wide growth parameter describing overall economic expansion. This definition allows inflation to be expressed in terms of the time evolution of this field.
In this framework, the purchasing power is represented by point $A$ in the field, and the price of goods by point $B$. The distance between both points is scaled by $x$. The availability of $B$ in the economic field (i.e., the presence of goods and services) is assumed to prevent deflationary collapse. The system is considered to be in an inflationary phase when the acceleration term satisfies $\ddot{x}>0$, and in a deflationary phase when $\ddot{x}<0$. An increase in the $x$ scale under the condition $\ddot{x}>0$ indicates a decrease in purchasing power, thus characterizing inflationary behavior.
Consequently, the acceleration of the economic field can be viewed as a measure of inflationary pressure, where fiscal and monetary interventions act as external forces capable of modulating the field’s expansion rate.
In the literature, the relative rate of change in the price of goods is generally defined as
$C=\frac{\dot{p}}{p}$
R37 ; R38 ; R39 ; R40 . According to the present framework, an alternative definition of the inflation rate can be expressed as
$C=\frac{\dot{x}}{x}$
which shows how an economic system or equivalently, an inflation field evolves over time. In such a system, the establishment of a developed economic equilibrium depends on the system’s ability to convert production value, represented by $B$, into purchasing power at point $A$.
In an inflationary context, the conversion of movable or immovable assets (including technological goods) into monetary resources often occurs through export activities or external resource inflows. These represent non-adiabatic processes within the global economic environment. On the other hand, it can be observed that the accessibility of $A$ to $B$ decreases continuously, since $C$ tends to decline over time under inflationary conditions. This behavior can, for instance, be modeled as
$x\simeq t^{h}$
where a decrease in $C$ occurs at a rate
$\frac{h}{t}$
when $h>1$. Therefore, the economic system exhibits a dynamic structure characterized by the scaling behavior of $x$.
Figure 1: Conceptual representation of a subsystem within the economic system $E$. Points AA and $B$ indicate two representative positions in the system, scaled by $x$. The element at AA corresponds to purchasing power, while $B$ represents the value of goods or assets. During inflationary expansion, systemic forces act outward, reflecting a general increase in the scale of economic transactions. Fiscal and monetary interventions function as regulatory mechanisms that help maintain balance between purchasing power and asset prices within this dynamic structure.
The energy density and pressure of this homogeneous and dynamic economic field are given as follows R41 ; R42 ,
$\rho=\frac{\dot{\varphi}^{2}}{2}+V(\varphi),\,\,\,\,\,\,\,\,\,\,p=\frac{\dot{\varphi}^{2}}{2}-V(\varphi)$
Here, the kinetic term $\frac{\dot{\varphi}^{2}}{2}$ corresponds to the purchasing power at point $A$, while the potential energy $V(\varphi)$ represents the accumulated value of movable and immovable goods at point $B$. A decrease in $C$ reflects reduced accessibility between purchasing power and goods prices. Inflation occurs when
$\frac{\dot{\varphi}^{2}}{2}<V(\varphi)$
whereas an economic equilibrium is achieved when
$\frac{\dot{\varphi}^{2}}{2}\simeq V(\varphi)$
The former corresponds to a phase of rapid expansion in the $x$-scale, characterized by high prices and low purchasing power, while the latter represents a stable, slowly expanding economy typical of developed systems. Therefore, the first case produces negative pressure associated with inflation (field expansion), whereas the second results in positive pressure reflecting purchasing power and economic stabilization.
A state parameter can be defined as
$\gamma=\frac{p_{\varphi}}{\rho_{\varphi}}$
which denotes the general state of the system. When $\gamma<0$, the system is in an inflationary (expanding) phase, while a positive $\gamma$ indicates a non-inflationary or decelerating condition.
Therefore, the analytical framework developed here allows inflation to be interpreted as a curvature-like expansion within an abstract economic field. This analogy highlights the role of fiscal and monetary interventions as control parameters that regulate the dynamic balance between purchasing power and asset values. While the proposed model remains phenomenological, it provides a conceptual bridge connecting macroeconomic behavior with the systemic dynamics of complex evolving structures observed across both physical and social domains.
We start with the Raychaudhuri equationR32 ; R33 ; R34 ; R35 , which describes the motion of localized units (densities) at points AA and $B$ defined in the field, and describes the motion between two points. For a homogeneous distributed dynamic system, this equation can be written as follows:
$\frac{d\theta}{d\tau}=-\frac{1}{3}\theta^{2}-4\pi(\rho_{\varphi}+3p_{\varphi}).$ (1)
where $\theta=\frac{\dot{x}}{x}$. Therefore, it can be written the equation,
$\frac{-3\ddot{x}}{x}=4\pi(\rho_{\varphi}+3p_{\varphi}).$ (2)
However, to achieve the continuity equation of the perfect fluid we make use of the first law of thermodynamics,
$\delta Q=dE+pdV=TdS,$ (3)
where $E$ is internal energy of subsystem (amount of money in the system), $V$ is the volume of system (the amount of goods in system), $p$ is the pressure of system (the indicator of purchasing power in the system) and $\delta Q$ is heat follow (money transfer that occurred between the subsystems or systems without buying and selling goods). Hence, since a macro-level observer will observe the energy from these quantities as a density dispersed in the volume in the subsystem, under the approximation $dE\simeq dm\simeq Vd\rho+\rho dV$, the first law is read as follows,
$\dot{\rho}+3C(\rho_{\varphi}+p_{\varphi})=\frac{\delta Q}{Vdt},$ (4)
This is the conservation equation of the economic system, which states that the quantities of the system are conserved during the economic cycle with time. Two solutions of equation (5) can be investigated for adiabatic ($\delta Q\simeq 0$) and non-adiabatic ($\delta Q\neq 0$) processes. The adiabatic case where there is no money transfer at the borders of the system means that there is no money transfer economically to a country without buying or selling goods. The other non-adiabatic process means there is a flow of money to or from country borders. In the next two sections, we will discuss the solutions of these two processes from an economic point of view.
II.1 Adiabatic process
The continuity equation for the adiabatic process takes the following form,
$\dot{\rho}+3C(\rho_{\varphi}+p_{\varphi})=0,$ (5)
and whose solution is as follows
$\rho_{\varphi}=\rho_{0}a^{-3(1+\gamma)}.$ (6)
Here, $\rho_{0}$ is an integral constant. Then, the eq. (5) takes the form,
$3\ddot{x}=4\pi\rho_{0}x^{-3(1+\gamma)+1}(1+3\gamma),$ (7)
So, we proceed by finding some solutions for the eq. (9) to investigate inflation. Using the modulating parameter,$w$, we can write $\dot{\varphi}^{2}(1-\gamma)=2(1+\gamma)V(\varphi)$. This equation tells us that there exists a boundary in an economic system at the point $\gamma\neq 1$ because the case $V(\varphi)=0$ representing movable or immovable property is not considered in an economic system. But, it is possible in the case $\gamma\sim-1$, which corresponds to a hyperinflation. This is a case where the monetary purchasing power represented by kinetic energy, $\dot{\varphi}^{2}$, is almost nonexistent or capital becomes almost worthless. Therefore, the solutions close to $\gamma\sim-1$ produce exponential hyperinflation solutions,
$x\simeq\exp(\sqrt{\frac{8\pi\rho_{0}}{3}}t)C_{1}+\exp(-\sqrt{\frac{8\pi\rho_{0}}{3}}t)C_{2},$ (8)
where $C_{1}$ and $C_{2}$ are integral constants. The first solution, which shows the increase in the distance scale between the purchasing power and the price of the good, is considered. In other words, an increase in scale $x$ means that there is a decreasing effect on purchasing power due to $\gamma\sim-1$ ($\frac{\dot{\varphi}^{2}}{2}\ll V(\varphi)$). Therefore, we can set $C_{2}$ to zero, where the solution $x\simeq\exp(\sqrt{\frac{8\pi\rho_{0}}{3}}t)C_{1}$ survives. This appears in Figure 1 with $C_{1}=1$.
Figure 2: The relationship between the scale variable $x$ and time $t$. The exponential increase in $x$ represents the hyperinflationary phase of the system. In this regime, purchasing power, interpreted as the system’s kinetic component, becomes almost negligible, while the rapid increase in prices is driven by accumulated asset values and capital concentration, corresponding to the potential component of the system.
The high potential indicates high prices of material or financial resources that have not yet been converted into monetary resources, that is, into the system’s kinetic component. An exponential increase in the scale variable $x$ indicates a loss of purchasing power. This unstable situation can be mitigated by reducing demand for goods, which in turn leads to a decline in potential, that is, a decrease in prices.
In an adiabatic system meaning no external monetary inflow or outflow the potential (price level) can be reduced either by increasing production or by lowering demand. It should be emphasized that artificially forcing the system into a non-adiabatic state is not an effective solution. Although a temporary monetary inflow (for example, capital entering the system through interest-based borrowing) can increase the kinetic component (purchasing power) in the short term, it ultimately leads to stronger inflation in the long run. This is reflected in Figure 2, where the curve peaks over time.
Furthermore, temporary external monetary inputs distort the homogeneity of the system by introducing artificial kinetic energy, which eventually drains resources through interest payments and creates structural imbalance. Monetary inflows enrich one segment while exploiting another, leading to an inhomogeneous field where purchasing power continues to erode. Hence, sustained inflation persists, especially for units whose kinetic energy (purchasing capacity) declines more rapidly.
Under adiabatic conditions, the most realistic path toward restoring equilibrium involves reducing demand encouraging economic savings and consumption restraint. This drives sellers to lower prices to convert their assets into money, which is a key mechanism for reducing inflation. In addition, increasing the availability of natural or productive resources decreases potential (prices) and enhances monetary valuation, moving the system toward stability.
However, wasting or underutilizing resources contradicts the policy direction of minimizing consumption. Therefore, satisfaction with existing goods either voluntarily or as a result of policy pressure becomes a natural adjustment mechanism.
The rate of inflation in this regime is given by $C=\rho_{0}$. If $\rho_{0}$ is fixed at a large value, inflation remains nearly constant over time. Thus, under rational policy frameworks, the rate of change $C$ should take a time-dependent form R42 to allow for gradual adjustment. In developed economies, the system tends to reach an equilibrium where purchasing power equals potential, $\frac{\dot{\varphi}^{2}}{2}\simeq V(\varphi)$, corresponding to $\gamma\simeq 0$. However, the solution of Eq. (10) near this point ($\gamma\simeq 0$) is an empty set.
Therefore, although $K=V$ holds approximately in adiabatic systems, under rational fiscal policies–such as demand reduction through savings–the system evolves toward a linear behavior at $\gamma\simeq-\frac{1}{3}$, where pressure gradually falls:
$x\sim C_{1}+tC_{2},$ (9)
Taking the real part ($C_{1}=0$), we obtain $x\sim tC_{2}$. The ratio parameter takes the form $C=\frac{1}{t}$, showing a decreasing trend over time. As a result, fiscal restraint policies that are maintained over long periods can effectively lower inflation.
If $t$ is measured in years, this model suggests that inflation could decline to single-digit levels over a period longer than ten years. As will be discussed in the following section, when moderate inter-system trade is allowed, an economic equilibrium point may be reached at $K=V$, and inflation may decline more rapidly within less than a decade even without strict fiscal constraints.
II.2 Non-adiabatic process
We shall use the entropic force definition to reveal how the term on the right-hand side of eq. (7). The entropic force is defined as follows,
$F_{ent.}=-\frac{TdS}{dx},$ (10)
where for a non-adiabatic process we have $TdS=\delta Q\neq 0$, so the eq. (7) can be written as follows,
$\frac{d\rho}{\rho}+3(1+\gamma)\frac{dx}{x}=-\frac{F_{ent.}dx}{E}.$ (11)
Herein, due to $E\simeq\rho V\simeq F_{ent.}x$, where recall that $\rho=kinetic+potential$, we can write
$\frac{d\rho}{\rho}+3(1+\gamma)\frac{dx}{x}=-\frac{dx}{x}.$ (12)
The solution of this equation is as follows,
$\rho=\rho_{0}x^{-3(1+\gamma)-1},$ (13)
with an integral constant $\rho_{0}$. Thus, the eq. (5) takes the form,
$3\ddot{x}=4\pi\rho_{0}(1+3\gamma)x^{-3(1+\gamma)},$ (14)
Hyperinflation solution $\gamma\simeq-1$ is
$x=\frac{4}{3}\pi\rho_{0}t^{2}+C_{1}+C_{2}t,$ (15)
where we set $C_{1}=C_{2}=0$, so an increase in $x$ is parabolic, indicating very low economic purchasing power.
Figure 3: Figure showing the increase of the x scale in the (0-8) time interval.
The proportional parameter of this type of inflationary field behaves as $C=\frac{2}{t}$. Even if the inflation parameter $C$, which starts in a zero time instance, returns to its natural appearance after 20 years, it reveals a deep deformation situation that ruins the field due to the size of the $x$ scale. Therefore, in such an inflationary field, it is necessary to intervene in the system with reasonable fiscal policies and to pass the inflationary phase to a new phase. In this way, a stable course can be followed by keeping the economy under control without any deformation.
Therefore, when we apply the $K=V$ or $w\simeq 0$ approximation that we discussed for economic equilibrium to the equation, we get the following equation.
$x=\pm\left(\frac{-4\pi\rho_{0}+3t^{2}C_{1}^{2}C_{2}+6tC_{1}^{2}C_{2}+3C_{1}^{2}C_{2}^{2}}{3C_{1}}\right)^{\frac{1}{2}}.$ (16)
Note that the solutions (17) and (18) converge at about $t=4$. We can manipulate the value of $C_{1}$ based on any value of $C_{2}$. For example, the value graph at $C_{1}=0.4$ for $C_{2}=1$ in the wide selected range of x=(415)x=(4-15) is given in Figure 1. If $C_{2}=2$ is chosen, $C_{1}=0.34$ produces the same graph. This means that inflation entered a declining phase (from $t\sim 4.1$ to about 11) and progressed in a developed economy phase after $t\sim 11$.
Figure 4: The figure shows the behavior of the x scale in the time interval (4-15). Here, $C_{1}=0.4$ is chosen.
We can again read the change in the inflation rate from the parameter $C=\frac{\dot{x}}{x}$. If we consider $t$ in terms of years, the economy with the traces of hyperinflation between $t=4.1$ and $t=5$, and a rate that decreases to 10 percent in approximately 7 years between $t=4.1$ and $t\simeq 11$ is obtained. After the approximate value of $t\simeq 11$, a single-digit inflation value may appear. This indicates the economic equilibrium represented by $K=V$ for the economic system. Let’s note that it takes about 10 years for it to reach equilibrium from a state of almost no purchasing power during the hyperinflation period and that it takes place with the external heat transfer to the system (occurring money transfer without buying-selling cases). Let’s remember that the adiabatic process did not produce the $K=V$ solution. In such processes, we emphasized that inflation can be avoided by reducing the demand for purchasing goods.
Country Inflation period
Bolivia 1970–1985
Peru 1970–1990
Israel 1970–1985
Brazil 1970–1994
Nicaragua 1985–1991
Hungary 1946 Jan–1946 Jul
Germany 1922 Nov–1923 Oct
Table 1: Inflation periods in certain year intervals of some countries R3 .
With the decrease in demand for the density units (energy, i.e., money) that describe the system, the kinetic energy or purchasing power will increase over time. Another possible scenario is political intervention aimed at reducing demand for the purchase of goods. Such policies reduce purchasing demand gradually, leading consumers to limit their needs, which increases kinetic energy over time and thus enhances purchasing power.
However, in the non-adiabatic process, external monetary investments or the presence of rich resources (potential assets convertible into money, i.e., kinetic energy) within the system will steer the economy toward balance over time, as our solutions have demonstrated.
Table 1 shows some countries that experienced hyperinflationary periods. Considering these intervals, it is observed that inflation typically persists for about 10 years, consistent with our theoretical calculations.
For example, under reasonable financial policies, the economic system in Turkey, which entered a hyperinflationary phase starting in 2017, may be expected to complete its natural course after 7 years and reduce inflation to single digits after approximately 11 years (around 2028).
III Closing remarks
There exist many symmetries in the universe (e.g. spatial and temporal translation symmetries, parity, and charge symmetries, etc.). Especially in physical science, symetry describes that when any observable quantity is considered, some properties of the system do not change under certain transformations (for example, Noether symmetries). In other words, symmetry tells us that the internal property of a physical system related to physics or mathematics does not change with the change of some factors. In social science, which is also a part of the universe, evaluating the welfare of society, the interaction of the units that make up the social system (including political interactions) and its development in terms of its compatibility with the laws of theoretical physics, can offer a realistic perspective. On the other hand, the physical development of a plant or human, the growth of the circle of ideas, or the expansion of the intellectual sphere of influence (just as the expansion of the universe) refer to observable symmetries. In this study, which we handled mathematically according to this point of view, which we call ”developmental symmetry” in the universe, we discuss inflation as an ever-expanding system (similar to the dynamic structure in developed economies) with its scalar field representation. In this direction, we investigate the economics of a system from an inflationary perspective within the framework of the Raychaudhuri equation. We have obtained a continuity equation based on the first law of thermodynamics. For the adiabatic and non-adiabatic processes of this equation, we obtain some solutions corresponding to the inflationary case for the system. As can be seen in Figure 1, the main point we consider in the study is the purchasing power, which is described by point $A$, and the purchased goods, which is identified by point $B$. It should be noted that in a developed economic system under consideration, the $x$-scale can never reach a value of zero. Finding an $x$ value in developed economies is important for the continuity and development of the economic system. Therefore, it is inevitable for a dynamic system (developed economic system) to have a constant value ($\frac{\dot{\phi}^{2}}{2}\simeq V(\phi)$) of purchasing power and price of goods for an exchange of approximately $x$ value to take place. An increase in the $x$ scale means that the availability of the good decreases, that is, the purchasing power ($\frac{\dot{\phi}^{2}}{2}$) decreases or the price of the good $V(\phi)$ is high. We represented this situation with $\frac{\dot{\phi}^{2}}{2}\ll V(\phi)$, therefore, solutions in which hyperinflation occurs in the case of $\frac{\dot{\phi}^{2}}{2}\ll V(\phi)$. As emphasized, the equilibrium point is the case of $\frac{\dot{\phi}^{2}}{2}\simeq V(\phi)$, which corresponds to a deflationary area that reduces the rate of hyperinflation. Since there is no energy (monetary) flow with external systems for an adiabatic economic system, it has been revealed that it can get rid of inflation in more than 10 years after it evolves into an inflation area of $\gamma=-\frac{1}{3}$ with reasonable financial policies within the system. Also considering the duration of reasonable policies, it will probably take more than 20 years to get out of the inflationary field. For the more realistic non-adiabatic process (the system with energy-money flow), we obtained mathematical results that are quite compatible with the inflation processes seen in Table 1. This analogy between thermodynamic balance and economic equilibrium also reflects the collective behavioral tendencies of economic agents–such as consumption habits, trust in fiscal policy, and market expectations–that determine how energy (money) circulates within a society.
On the other hand, the main reason why some countries in Table 1 left the hyperinflation area in less than 11 years can be shown as the inflow of money into the country. For example, the cash flow from the international monetary fund to the country arising from production, natural resources, technological infrastructure and the sound economic policies created as a result can be shown. The reasons for the inflation that continues for more than 11 years on average may be wrong fiscal policies and management of the system (country) as an adiabatic system. In our current study, the 11-year average period of the inflation process is obtained within the framework of ”developmental symmetry” for non-adiabatic systems. After an average of 11 years of hyperinflation, the economic system is evolving into a stable state (single-digit percentages). As we have emphasized, one of the most effective ways to stop hyperinflation in such an economic system created identically (symmetrically) to mathematical tools is to ensure production in the system. Whether such a system is adiabatic or not (whether there is energy-money flow or not) may no longer be important. The use of natural resources with the right policies or the production of technological parameters that may meet the needs can lead to a reasonable inflation rate without the flow of money with interest to the inflation area where purchasing power drops to almost zero. In another separate study, it can be discussed whether inflation can be cut or not by finding the creation of the scalar field density in the field. Because density is a quantity that includes kinetics and potential due to $E\sim\rho$. Units of density created in the field by the political ground controlling the field represents the unity of purchasing power (kinetic energy, point $A$ in Figure 1) and movable immovable property (potential energy, point $B$ in Figure 1). Therefore, it is clear that this unity emerging everywhere in the field will have a decreasing effect on inflation. This highlights that maintaining the balance between purchasing power and asset value throughout the economic field is crucial for mitigating hyperinflation and ensuring sustainable economic growth.
IV Data Availability Statement
No Data associated in the manuscript